OpenStudy (anonymous):

square root of 7-x=19

OpenStudy (anonymous):

what is the square root of -12

OpenStudy (anonymous):

sqrt(7-x)=19 7-x=(19)^2 7-x=361 -x=354 x=-354

OpenStudy (anonymous):

3.46410162